OpenStudy (anonymous):

Consider the set of \(n\times n\) matrices \(M_n\) where the entry in row \(i\) and column \(j\) is defined by\[m_{i,j}=\begin{cases}0&\text{for }i=j\\[1ex]1&\text{otherwise}\end{cases}\]so for instance,\[M_2=\begin{pmatrix}0&1\\[1ex]1&0\end{pmatrix}\quad\text{and}\quad M_3=\begin{pmatrix}0&1&1\\[1ex]1&0&1\\[1ex]1&1&0\end{pmatrix}\]and so on. Show that for positive integers \(n\ge1\), \[\det M_n=(-1)^{n-1}(n-1)\]

OpenStudy (anonymous):

I stumbled across this pattern while playing around with a system of recurrence relations that correspond to the matrices \(M_n\) for \(n\ge2\) (but it's true for \(n=1\)). I've been trying to prove this by induction, but the inductive step is tricky to me. Manually crunching the numbers via cofactor expansions gets messy quickly. The pattern holds, ie. \(\det M_2=-1\) and \(\det M_3=2\) and so on, and I've verified with Mathematica for up to \(n=100\). Aside - the system of recurrence relations I mentioned. I've been looking at the system\[\begin{cases}a_0=A\\[1ex]b_0=B\\[1ex]a_{n+1}=b_n\\[1ex]b_{n+1}=a_n\end{cases}~~\iff~~\begin{pmatrix}a_{n+1}\\[1ex]b_{n+1}\end{pmatrix}=\begin{pmatrix}0&1\\[1ex]1&0\end{pmatrix}\begin{pmatrix}a_n\\[1ex]b_n\end{pmatrix}\]and the extension to \(n\) equations\[\begin{pmatrix}a_{1,k+1}\\[1ex]\vdots\\[1ex]a_{n,k+1}\end{pmatrix}=M_n\begin{pmatrix}a_{1,k}\\[1ex]\vdots\\a_{n,k}\end{pmatrix}\]and it turns out that the determinant of each \(M_n\) seems to play a role in one of the coefficients of the generating functions for the sequences \(a_i\).

OpenStudy (anonymous):

The cofactor expansions for small \(n\) gave me a basic pattern of being able to reduce the determinant of the initial matrix \(\det M_n\) into significantly simpler determinants of \(2\times2\) matrices, and these determinants are exclusively \(1\), \(0\), or \(-1\). I don't think this helps in a proof by induction, though, but I'm not entirely convinced. Unless there's some way to inductively (or otherwise) prove that \[\det M_n=\sum (\det m_a+\det m_b+\det m_c)\]where \(m_a,m_b,m_c\) are all \(2\times2\) matrices with determinants \(1\), \(0\), \(-1\) respectively.

ganeshie8 (ganeshie8):

alternatively, elimination may work nicely

OpenStudy (anonymous):

I've also considered writing \(M_n\) as \(C_n-I_n\), where \(C_n\) is a constant matrix composed entirely of \(1\)s, but alas, \(\det(C_n-I_n)\neq\det C_n-\det I_n\).

ganeshie8 (ganeshie8):

1) \(R_1 + R_2+\cdots + R_n\) makes all the elements in the first row equal to \((n-1)\) 2) pull \(n-1\) out from first row 3) do \(R_i - R_1\) for other rows to get a upper triangular matrix with \(-1\) in the diagonal in all the below \(n-1\) rows 4) multiply the entries in diagonal to get determinant

OpenStudy (anonymous):

Oh good call on the triangulizing (?)

ganeshie8 (ganeshie8):

that looks more like a construction based proof...im still thinking of induction.. .

ganeshie8 (ganeshie8):

hey may i see the generating function of \(\{a_n\}\)

OpenStudy (anonymous):

Still, a useful technique! A cofactor expansion of \(M_{k+1}\) gives \[\det M_k+\sum(\text{other matrices})=(-1)^{k-1}(k-1)+\cdots\]Cheating a bit, and "knowing" the RHS should add up to \((-1)^kk\), we get \[\sum(\text{other matrices})=(-1)^k(2k-1)\]but whether that gives any useful results, I'm not sure.

OpenStudy (anonymous):

Here are the GFs for \(a_n\) and \(b_n\) in the case of the \(2\) recurrence system. I got \[a(x)=\color{red}{\frac{A+Bx}{1-x^2}}=\sum_{k\ge0}\left(\frac{A+B}{2}+(-1)^n\frac{A-B}{2}\right)x^n\\ b(x)=\color{blue}{\frac{B+Ax}{1-x^2}}=\sum_{k\ge0}\left(\frac{A+B}{2}+(-1)^n\frac{B-A}{2}\right)x^n\](possibly a typo somewhere, I wouldn't be surprised)

OpenStudy (anonymous):

I've only solved the double recurrence so far, and I have the GFs for the triple recurrence, but I've already made a basic conjecture that \(\det M_n\) gives the coefficient of the highest power term in the denominator (possibly multiplied by a power of \(-1\))

OpenStudy (anonymous):

Here's the set of GFs for the triple recurrence: \[\begin{cases}a_0=A;\quad b_0=B;\quad c_0=C\\[1ex]a_{n+1}=b_n+c_n\\[1ex]b_{n+1}=a_n+c_n\\[1ex]c_{n+1}=a_n+b_n\end{cases}\]\[a(x)=\frac{Ax-(A-B-C)x^2}{1-x-2x^2}\\ b(x)=\frac{Bx-(B-A-C)x^2}{1-x-2x^2}\\ c(x)=\frac{Cx-(C-A-B)x^2}{1-x-2x^2}\]I wish I had the time and patience to work with the quadruple recurrence just to see if there's any merit to the conjecture about the leading coefficient.

OpenStudy (anonymous):

Interesting observation (though it's actually not completely surprising now that I think about it), for the triple recurrence, the \(k\)th term of the sequence \(\{a_n,b_n,c_n\}\) is given by \({M_3}^k\begin{pmatrix}A\\B\\C\end{pmatrix}\). It's probably true then that the \(k\)th term of \(\{a_{1,n},\ldots\,a_{k,j}\}\) is given by \({M_n}^k\begin{pmatrix}A_1\\\vdots\\A_j\end{pmatrix}\). Pretty neat.

OpenStudy (kainui):

I came here to try to figure out and say what ganeshie already did, but I am still figuring stuff out and interested in your recurrence problems more this looks fun.

OpenStudy (kainui):

Interesting, it appears that my linear algebra workbook has this fun example of showing that: \[\det[M_5 -4I]=0\] Which probably generalizes to this (if you don't already know that from earlier I can't say) \[\det[M_n +(n-1)I]=0\]

OpenStudy (anonymous):

\(M_n\)s the same matrix, right?

OpenStudy (kainui):

Yes, actually they have the same exact problem as you're asking, verbatim here it is: Show that the n-square determinant \[\begin{vmatrix} 0 &1 &1 &\cdots &1 \\1 &0 &1 &\cdots &1 \\1 &1 &0 &\cdots &1 \\\cdots &\cdots &\cdots &\cdots &\cdots \\\cdots &\cdots &\cdots &\cdots &\cdots \\1 &1 &1 &1 &0 \end{vmatrix} = (n-1)\begin{vmatrix} 0 &1 &1 &\cdots &1 \\1 &0 &1 &\cdots &1 \\1 &1 &0 &\cdots &1 \\\cdots &\cdots &\cdots &\cdots &\cdots \\\cdots &\cdots &\cdots &\cdots &\cdots \\1 &1 &1 &1 &1 \end{vmatrix} = (-1)^n (n-1)\]

OpenStudy (anonymous):

And here I thought I was being so creative :P

OpenStudy (kainui):

Here's another problem in the book! \[\begin{vmatrix} a+b &a &a &\cdots &a \\a &a+b &a &\cdots &a \\\cdots &\cdots &\cdots &\cdots &\cdots \\a &a &a &\cdots &a+b \end{vmatrix} = b^{n-1}(na+b)\]

OpenStudy (anonymous):

I've also been thinking about another set of matrices with this pattern: \[{M'}_2=M_2\quad\quad{M'}_3=\begin{pmatrix}0&1&2\\[1ex]1&0&1\\[1ex]2&1&0\end{pmatrix}\quad\quad{M'}_4=\begin{pmatrix}0&1&2&3\\[1ex]1&0&1&2\\[1ex]2&1&0&1\\[1ex]3&2&1&0\end{pmatrix}\]and so on. These also have a cool alternating pattern with their determinants: \(-1,4,-12,32,\ldots\).

OpenStudy (anonymous):

So basically the entry in row \(i\), column \(j\) is \(|i-j|\).

OpenStudy (kainui):

Oooh ok these look cool, they're not just symmetric they're centrosymmetric! https://en.wikipedia.org/wiki/Centrosymmetric_matrix The exchange matrix J is defined as the anti-diagonal (top right to bottom left) version of the identity matrix, then your matrices satisfy the centrosymmetric equation: AJ=JA just like a symmetric matrix satisfies \(A=A^\top\) I don't know if that would help/be interest to you, also this reminds me of the pascal triangle matrices \[UL=P\] So upper triangular pascal matrix times lower triangular pascal matrix gives the full pascal matrix... (might have to put in a third column and row to see the pattern better: \[\begin{bmatrix} 1 &0 \\1 &1 \end{bmatrix}\begin{bmatrix} 1 &1 \\0 &1 \end{bmatrix}=\begin{bmatrix} 1 &1 \\1 &2 \end{bmatrix} \] Some related problems or ideas for inspiration for now while I think of other stuff.

OpenStudy (anonymous):

Haha I was wondering if these types of matrices had a name. For my original \(M_n\) I was thinking something like antidiagonal or "pseudodiagonal", but I found the former was already taken.

OpenStudy (kainui):

Yeah, there's Persymmetric, Centrosymmetric, Bisymmetric, and Toeplitz matrices which all seem to be pretty closely related, I've looked into this before just from playing around with some permutations since the exchange matrix J had interested me when I tried to solve some problem a while back.

OpenStudy (kainui):

So these \(M_n'\) matrices you've introduced here are interesting, all I can tell so far is that their determinant is always divisible by (n-1) due to the fact that adding the first column to the last column produces a column with all 1's in it multiplied by (n-1).

OpenStudy (anonymous):

Hmm, I wonder how that result might change if we mirror the matrices across their centers (such that \({M'}_2\) becomes \(I_2\), for instance).

OpenStudy (kainui):

Oh could you explain that a little more what you mean by mirroring, I'm not sure I follow.

OpenStudy (kainui):

Although probably far from the beaten path, I still find this to be kind of interesting. You can represent every matrix as the sum of a centrosymmetric and skew-centrosymmetric matrix: $$A = \frac{A+JAJ}{2}+\frac{A-JAJ}{2} = B+C$$ \(BJ=JB\) (hehe) and \(CJ=-JC\) are the relations for centrosymmetric and skew-centrosymmetric

OpenStudy (anonymous):

By mirroring (maybe there's an existing term that's more technical) I mean some transformation, let's call it \(\mathrm{Mirv}\), that gives \[\mathrm{Mirv}\begin{pmatrix}1&2&3\\[1ex]4&5&6\\[1ex]7&8&9\end{pmatrix}=\begin{pmatrix}3&2&1\\[1ex]6&5&4\\[1ex]9&8&7\end{pmatrix}\]or another transformation \(\mathrm{Mirh}\) which gives the transpose of a "mirv'd" matrix. (v for reflection along some vertical axis, h for horizontal)